How fast must a 58 kg mountaineer climb a hill that has a slope of 16° to come up with a power output of 1.0 hp?


1
Expert's answer
2022-01-17T09:58:50-0500

Given:

"m=58\:\rm kg"

"\theta=16^\circ"

"P=1.0\rm \: hp=746\: W"


The power is related to the speed and force by an equation

"P=Fv=mg\sin\theta v"

Hence, the speed of a mountaineer

"v=\frac{P}{mg\sin\theta}\\
=\frac{746}{58*9.8\sin16^\circ}=4.8\:\rm m/s"